OpenStudy (anonymous):

Write the simplest polynomial function with zeros 4+i and sqrt of 2? help?!

OpenStudy (aum):

x = 4+i is one root x - 4 = i square both sides: (x-4)^2 = i^2 x^2 - 8x + 16 = -1 x^2 - 8x + 17 = 0 x = sqrt(2) (x - sqrt(2)) is a factor. Assuming irrational coefficients are okay, just multiply (x-sqrt(2)) and (x^2 - 8x + 17) and simplify.